HUBBLE SPACE TELESCOPE DAILY REPORT #1290 PERIOD COVERED: 0000Z Daily Status Report as of 355/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5563 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1: UV /OPT STD.) The WF/PC-2 was used to make photometric calibration images of ultraviolet standard star GRW+70D582 and monitor the photometric stability and quantum efficiency of the WF/PC-2. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. There was a tape recorder track change while a observation was writing to tape; this resulted in a data loss of one group. No other anomalies were noted. 1.3 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 1.4 Completed FOS/RD 5391 (Mapping the Hydroxyl Distribution in the Saturn Magnetosphere: Cycle 4) The Faint Object Spectrograph (Red Side) was used to make observations of regions near the planet Saturn (SATH1) to study the distribution of hydroxyl radical (OH) in its magnetosphere. Earlier investigations with HST have found the amount of OH to be unexpectedly high. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed GHRS/2 5600 (GHRS Side 2 Wavelength Monitor (1st Order Gratings): Cycle 4) The High Resolution Spectrograph (Side Two) was used to make wavelength and resolution calibration observations of its first order gratings using the internal calibration lamp. The observations were completed as planned, and no anomalies were reported. 1.6 Completed FOC/96 5175 (Physics of Asteroids: Cycle 4) The Faint Object Camera (F/96) was used to make observations of the asteroid Vesta to look for surface features and to determine the shape and diameter of the asteroid. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FGS 5657 (Duplicity Among Hyades Stars: Cycle 4 High) The FGSs (TRANS-Mode) were used to make astrometric observations of star VA709 located in the Hyades cluster in an attempt to find a close binary. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program: Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 5401 (Luminosities and Morphologies of a Complete 3CR Sample of Radio Galaxies) The WF/PC-2 was used to obtain images of radio galaxy 3C280, which has a Z near 1.0, and also is a quasar. These observations will be used to determine and compare luminosities, and morphologies in the optical and infrared. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Notes: HST operations continue to be supported via WSGT as the STGT is now re-scheduled to recover from its power problems at 356/0630Z. The modified "H" format was uplinked yesterday and activated for use at 354/1300Z. This format was utilized for the RMGA full calibration which was nominally completed at 354/1734Z. The Safemode analysts commanded the PSEA and monitored the telemetry without any difficulty. The calibration software, which had been previously modified via a workaround, worked flawlessly. The RMGA drift rate bias continues show a slight change for RMGA 2 and no change for RMGAs 1 and 3. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-